the measure of the angles of a triangle are in the ratio of 2:3:7. Find the measure of each angle.

Found 2 solutions by Jk22, mananth:
Answer by Jk22(389)

2x+3x+7x=180=12x, 2x=30, x=15, the angles are :
2x=30, 3x=45, 7x=105

Answer by mananth(16949)

the ratios are 2:3:7
Total of the ratio = 12
..
first angle = 2/12 *180
=30
..
second angle = 3/12 * 180
=45
..
third angle = 7/12 *180=105
...
30+45+105 = 180
...
m.ananth@hotmail.ca